Shift report 20011028 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Oct 28 2001 - 00:19:29 EDT

  • Next message: Apache: "Shift report 20011028 00:00-08:00"

    ------------------------------------------------------------
    Date: Sun Oct 28 00:19:29 2001, Shift: 16:00-24:00
    Supervisor: R.Wada, with: 
    Log book#10 pages 63 - 63
    
    Shift Summary:
    ------------------------------------------------------------
    1. After three times failures of ramping at 16.10, 17.15 and 18.15,
    machine went into Ramping efficiency study mode until 22:10. At 22:30
    finally they succeeded to store and cogg the beam. Data taking
    started at 23:10. stopped at 00:03.
    2. run 5538 TPC pedestal run(wromg run type. palser instead of pedestal.)
           5539 DVM
           5540 Pedestals
    
           5541 data run FS @3deg. w 1/2A, MRS @40deg. w 3A,
                 same as the previous data run.  triger 6 61k
    3. No new problem is noticed.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Oct 28 2001 - 00:20:32 EDT